OpenStudy (hellokitty17):

The following graph shows the functions f(x) and g(x): graph of function f of x equals x squared and graph of function g of x equals x squared plus 7 The function g(x) is obtained by adding _______ to f(x). (only input integers) Numerical Answers Expected!

OpenStudy (hellokitty17):

@ikram002p

OpenStudy (hellokitty17):

@mathmath333

OpenStudy (hellokitty17):

hey help me please

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} f(x)=x^2,\ g(x)=x^2+7\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

this is same as \(\large \color{black}{\begin{align} y=x^2,\ y=x^2+7\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (hellokitty17):

ok

OpenStudy (mathmath333):

graph of \(y=x^2\) is 

OpenStudy (hellokitty17):

yes and what would u add

OpenStudy (mathmath333):

graph of \(y=x^2+7\) is 

OpenStudy (hellokitty17):

so u would add 7

OpenStudy (mathmath333):

i think it easy u have to add \(7\)

OpenStudy (hellokitty17):

yup that is what I said